# Hyperelliptic four-manifolds defined by vector-colorings of simple polytopes

### v2

@@ Abstract.
The work was funded within the framework of the HSE University Basic Research Program
@@ 1. Introduction
Toric topology (see [BP15, DJ91]) assigns to each simple $n$-polytope with $m$ facets $F_{1},\dots,F_{m}$ the real moment-angle manifold
@@
by the facets $S^{n}_{k,\geqslant}\cap\{x_{i}=0\}$, $i=1,\dots,k$, and ([E24, Corollary]) to the complex given on
@@
by the facets $B^{n}_{k-1,\geqslant}\cap\{x_{s}=0\}$, $s=1,\dots,k-1$, and $B^{n}_{k-1,\geqslant}\cap\{x_{1}^{2}+\dots+x_{n}^{2}=1\}$. We denote the equivalence class of these complexes $\mathcal{C}(n,k)$. For $n=3$ the complex $\mathcal{C}(3,1)$ has a single facet $\partial P$, the complex $\mathcal{C}(3,1)$ corresponds a simple edge-cycle in $\partial P$, $\mathcal{C}(3,2)$ – to a theta-subgraph in $\partial P$ consisting of two disjoint vertices of $P$ and three disjoint (outside these vertices) simple edge-paths connecting them, and $\mathcal{C}(3,4)$ corresponds to a $K_{4}$-subgraph in $\partial P$ consisting of $4$ disjoint vertices of $P$ pairwise connected by a set of disjoint simple edge-paths.
@@
In dimension $n=3$ in the papers [M90, VM99M, VM99S2] A.D. Mednykh and A.Yu.Vesnin constructed examples of hyperelliptic $3$-manifolds in five of eight Thurston’s geometries: $\mathbb{S}^{3}$, $\mathbb{R}^{3}$, $\mathbb{L}^{3}$, $\mathbb{S}^{2}\times\mathbb{R}$, and $\mathbb{L}^{2}\times\mathbb{R}$. Each manifold was constructed using a Hamiltonian cycle, a Hamiltonian theta-subgraph or a Hamiltonian $K_{4}$-subgraph in a compact right-angled $3$-polytope, where a subgraph is Hamiltonian, if it contains all the vertices of $P$. In [E24, Theorem 11.5] it was proved that for a $3$-manifold $N(P,\Lambda)$ over a $3$-polytope hyperelliptic involutions in $\mathbb{Z}_{2}^{r}$ are in bijection with a Hamiltonian empty set, cycles, theta-subgraphs and $K_{4}$-subgraphs in $\mathcal{C}(P,\Lambda)$ inducing $\Lambda$, where for a Hamiltonian subgraph $\Gamma$ of the above types in the $1$-skeleton of $\mathcal{C}(P,c)$ there is a canonical way to construct a vector-coloring $\widetilde{\Lambda}_{\Gamma}$ of $P$ induced by $\Gamma$ such that $\mathcal{C}(P,c)=\mathcal{C}(P,\widetilde{\Lambda}_{\Gamma})$. Also in [E24, Theorem 11.7] the classification of vector-colorings with more than one hyperelliptic involutions was presented.
In this paper we generalize [E24, Theorem 11.5] to dimension $n=4$. For this purpose we introduce the notion of a Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C\subset\mathcal{C}(P,c)$ (Definition 4.32) and explain (Construction 4.38) how to build a canonical induced vector-coloring $\widetilde{\Lambda}_{C}$ of rank $k+1$ such that $\mathcal{C}(P,c)=\mathcal{C}(P,\widetilde{\Lambda}_{C})$ and $N(P,\widetilde{\Lambda}_{C})$ has a canonical hyperelliptic involution in $\mathbb{Z}_{2}^{k+1}$. Then for a vector-coloring $\Lambda$ of rank $r$ any Hamiltonian $\mathcal{C}(n,r-1)$-subcomplex $C\subset\mathcal{C}(P,\Lambda)$ inducing $\Lambda$ corresponds to a hyperelliptic involution $\tau_{C}\in\mathbb{Z}_{2}^{r}$. Our second main result (Theorem 4.39) is that in dimensions $n\leqslant 4$ this correspondence is a bijection.
@@
In Subsection 4.3 using the notion of a bipartite Hamiltonian subcomplex we reformulate (Corollary 4.20) the four color theorem for a simple $3$-polytope $P$ in terms of a disjoint union of circles containing all the vertices of $P$.
@@
  • In Proposition 4.29 and Corollary 4.30 we describe $k$-faces of a Hamiltoninan subcomplex $C\subset\mathcal{C}(P,c)$ as unions of $k$-faces of $\mathcal{C}(P,c)$ separated by (k1)(k-1)-faces of defining faces.
@@
  • In Corollary 4.33 we describe eplicitly the subgroup $H(C)\subset\mathbb{Z}_{2}^{m}$ corresponding to a $\mathcal{C}(n,k)$-subcomplex $C\subset\partial P$.
@@
  • In Construction 4.38 we show that any proper Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C\subset\mathcal{C}(P,c)$ induces a canonical vector-coloring $\widetilde{\Lambda}_{C}$ of rank $(k+1)$ such that $\mathcal{C}(P,c)=\mathcal{C}(P,\widetilde{\Lambda}_{C})$ and $N(P,\widetilde{\Lambda}_{C})$ is a hyperelliptic manifold with a canonical hyperelliptic involution in $\mathbb{Z}_{2}^{k+1}$ corresponding to $C$.
  • In Theorem 4.39 we formulate our second main result, namely that for $n\leqslant 4$ and a vector-coloring of rank $r$ of a simple $n$-polytope $P$ such that $N(P,\Lambda)$ is a closed manifold hyperelliptic involutions are in bijection with proper Hamiltonian $\mathcal{C}(n,r-1)$-subcomplexes $C\subset\mathcal{C}(P,\Lambda)$ inducing $\Lambda$.
@@
In Subsection 4.7 we study geometric structures on hyperelliptic manifolds $N(P,\Lambda)$. Our main tool is Construction 4.45 of a geometric hyperelliptic manifold using a proper Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C$ in the boundary of a compact right-angled polytope $P$ in some geometry $\mathbb{X}$. This construction is a direct generalization of the construction from [M90, VM99M, VM99S2] based on a Hamiltonian cycle, theta- or $K_{4}$-subgraph in the boundary of a right-angled $3$-polytope.
@@
  • In Corollary 4.49 we show that $\mathbb{I}^{n}$ does not admit Hamiltonian $\mathcal{C}(n,k)$-subcomplexes for $n\leqslant 4$. In particular, the geometry $\mathbb{R}^{n}$ does not admit Construction 4.45 for $n\geqslant 4$.
@@
  • In Corollaries 4.51 and 4.53 and Remark 4.52 we show that if $P$ admits a Hamiltonian $\mathcal{C}(n,n-1)$-subcomplex, then each defining face is a polytope with only even-gonal 22-faces.
@@
  • In Proposition 4.56 we show that if a $4$-polytope $P$ admits a proper Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then $P$ has at least one $2$-face with three or four edges. In particular, the geometry $\mathbb{L}^{4}$ does not admit Construction 4.45, as well as $\mathbb{L}^{n}$ for $n\geqslant 5$ since in the latter geometries there are no compact right-angled polytopes due to results by V.V. Nikulin [N82] (see [V17]).
@@
  • In Proposition 4.57 we show that if a $4$-polytope $P=Q\times I$ admits a proper Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then one of the defining faces is a triangle. In particular, the geometries $\mathbb{L}^{3}\times\mathbb{R}$ and $\mathbb{L}^{2}\times\mathbb{R}^{2}$ do not admit Construction 4.45.
@@
  • In Proposition 4.58 we show that $\Delta^{n}$ admits a Hamiltonian $\mathcal{C}(n,n)$-subcomplex. In particular, the geometry $\mathbb{S}^{n}$ admits Construction 4.45 for $n\geqslant 2$.
@@
  • In Proposition 4.61 we show that the prism $\Delta^{n-1}\times I$ admits Hamiltonian $\mathcal{C}(n,n)$ and $\mathcal{C}(n,n+1)$-subcomplexes for $n\geqslant 4$. In particular, the geometry $\mathbb{S}^{n-1}\times\mathbb{R}$ admits Construction 4.45 for $n\geqslant 3$.
@@
  • In Construction 4.62 we describe the operation of cutting off a face and show that the initial polytope defines a proper Hamiltonian subcomplex in the resulting polytope. As a result in Corollary 4.68 we show that the polytope $P^{n}=\Delta^{p}\times\Delta^{q}$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex for $p,q\geqslant 1$. In particular, the geometry $\mathbb{S}^{p}\times\mathbb{S}^{q}$, $p,q\geqslant 2$, admits Construction 4.45.
@@
  • In Proposition 4.70 we show that if $P^{n}$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex such that its defining faces $\{M_{1},\dots,M_{s}\}$ do not intersect a facet $F_{i}\simeq\Delta^{n-1}$, then for any $k\geqslant 1$ the polytope $\Delta^{k}\times P$ has a Hamiltonian $\mathcal{C}(n+k,n+k+1)$-subcomplex, and in Corollary 4.71 we show that for any $p,q\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times\Delta^{q}\times I$ has a Hamiltonian $C(n,n+1)$-subcomplex. In particular, the geometries $\mathbb{S}^{p}\times\mathbb{S}^{q}\times\mathbb{R}$, $p,q\geqslant 2$, and $\mathbb{S}^{p}\times\mathbb{R}^{2}$, $p\geqslant 2$, admit Construction 4.45.
@@
  • In Proposition 4.72 we show that if an $n$-polytope $P$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex $C$ such that its defining faces $\{M_{1},\dots,M_{s}\}$ do not intersect an (n2)(n-2)-face $G\subset P$, then $G\simeq\Delta^{n-2}$, and $P$ has a Hamiltonian $\mathcal{C}(n,n)$-subcomplex. Using this fact in Corollary 4.73 we show that $\Delta^{2}\times\Delta^{n-2}$ has a Hamiltonian $\mathcal{C}(n,n)$-subcomplex for $n\geqslant 3$, and in Corollary 4.74 we show that for $p\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times P_{k}$, where $P_{k}$ is a $k$-gon, $k\geqslant 3$, admits Hamiltonian $\mathcal{C}(n,n)$- and $\mathcal{C}(n,n+1)$-subcomplexes. In particular, the geometries $\mathbb{S}^{p}\times\mathbb{S}^{2}$, $\mathbb{S}^{p}\times\mathbb{R}^{2}$ and $\mathbb{S}^{p}\times\mathbb{L}^{2}$, $p\geqslant 2$, admit Construction 4.45.
@@
  • In Example 4.75 we show that the polytope $P_{5}\times P_{5}$ admits a Hamiltonian $\mathcal{C}(4,5)$-subcomplex. In particular, the geometry $\mathbb{L}^{2}\times\mathbb{L}^{2}$ admits Construction 4.45.
@@ 2. Rational homology $4$-spheres among $N(P,\Lambda)$
the Poincare-Lefschets duality $H_{k}(X,\partial X)\simeq H^{n-k}(X)$ and the universal coefficients formula. ∎
@@
[E24, Lemma 2.2] Let a point $p\in\partial P$ belong to exactly $l\geqslant 0$ facets $G_{i_{1}}$, $\dots$, $G_{i_{l}}$ of $\mathcal{C}(P,c)$. Then there is a piecewise linear homeomorphism $\varphi$ of a neighbourhood $U\subset P$ of $p$ such that $U\cap G_{j}=\varnothing$ for $j\notin\{i_{1},\dots,i_{l}\}$ to a neighbourhood $V\subset\mathbb{R}^{l}_{\geqslant}\times\mathbb{R}^{n-l}$ such that $\varphi(G_{j_{s}}\cap U)=V\cap\{y_{s}=0\}$, $s=1,\dots,l$.
@@
Let $P$ ne an $n$-polytope, and let $\omega_{1},\omega_{2}\subset[m]$ be subsets such that $\omega_{1}\cap\omega_{2}=\varnothing$, and $P_{\omega_{1}}$, $P_{\omega_{2}}$ are (n1)(n-1)-RHDRHD.
@@
Let $P$ ne an $n$-polytope, $n\leqslant 4$, and let $\omega_{1},\omega_{2}\subset[m]$ be subsets such that $\omega_{1}\cap\omega_{2}=\varnothing$, and $P_{\omega_{1}}$, $P_{\omega_{2}}$ are (n1)(n-1)-disks.

### v3

@@ Abstract.
The study has been funded within the framework of the HSE University Basic Research Program
@@ 1. Introduction
Toric topology (see [BP15, DJ91]) assigns to each simple convex $n$-polytope $P$ with $m$ facets $F_{1},\dots,F_{m}$ the real moment-angle manifold
@@
by the facets $S^{n}_{k,\geqslant}\cap\{x_{i}=0\}$, $i=1,\dots,k$, and ([E24, Corollary 2.11]) to the complex given on
@@
by the facets $B^{n}_{k-1,\geqslant}\cap\{x_{s}=0\}$, $s=1,\dots,k-1$, and $B^{n}_{k-1,\geqslant}\cap\{x_{1}^{2}+\dots+x_{n}^{2}=1\}$. We denote the equivalence class of these complexes $\mathcal{C}(n,k)$. For $n=3$ the complex $\mathcal{C}(3,1)$ has a single facet $\partial P$, the complex $\mathcal{C}(3,1)$ corresponds a simple edge-cycle in $\partial P$, $\mathcal{C}(3,2)$ – to a theta-subgraph in $\partial P$ consisting of two different vertices of $P$ and three disjoint (outside these vertices) simple edge-paths connecting them, and $\mathcal{C}(3,4)$ corresponds to a $K_{4}$-subgraph in $\partial P$ consisting of $4$ disjoint vertices of $P$ pairwise connected by a set of disjoint simple edge-paths.
@@
In dimension $n=3$ in the papers [M90, VM99M, VM99S2] A.D. Mednykh and A.Yu.Vesnin constructed examples of hyperelliptic $3$-manifolds in five of eight Thurston’s geometries: $\mathbb{S}^{3}$, $\mathbb{R}^{3}$, $\mathbb{L}^{3}$, $\mathbb{S}^{2}\times\mathbb{R}$, and $\mathbb{L}^{2}\times\mathbb{R}$. Each manifold was constructed using a Hamiltonian cycle, a Hamiltonian theta-subgraph or a Hamiltonian $K_{4}$-subgraph in the $1$-skeleton of a compact right-angled $3$-polytope, where a subgraph is Hamiltonian, if it contains all vertices of the polytope. In [E24, Theorem 11.5] it was proved that for a $3$-manifold $N(P,\Lambda)$ over a $3$-polytope $P$ hyperelliptic involutions in $\mathbb{Z}_{2}^{r}$ are in bijection with a Hamiltonian empty set, cycles, theta-subgraphs and $K_{4}$-subgraphs in the $1$-skeleton of $\mathcal{C}(P,\Lambda)$ inducing $\Lambda$, where for a Hamiltonian subgraph $\Gamma$ of the above types in the $1$-skeleton of $\mathcal{C}(P,c)$ there is a canonical way to construct a vector-coloring $\widetilde{\Lambda}_{\Gamma}$ of $P$ induced by $\Gamma$ such that $\mathcal{C}(P,c)=\mathcal{C}(P,\widetilde{\Lambda}_{\Gamma})$. Also in [E24, Theorem 11.7] the classification of vector-colorings with more than one hyperelliptic involutions was presented.
In this paper we generalize [E24, Theorem 11.5] to dimension $n=4$. For this purpose we introduce the notion of a Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C\subset\mathcal{C}(P,c)$ (Definition 4.32) and explain (Construction 4.39) how to build a canonical induced vector-coloring $\widetilde{\Lambda}_{C}$ of rank $k+1$ such that $\mathcal{C}(P,c)=\mathcal{C}(P,\widetilde{\Lambda}_{C})$ and $N(P,\widetilde{\Lambda}_{C})$ has a canonical hyperelliptic involution in $\mathbb{Z}_{2}^{k+1}$. Then for a vector-coloring $\Lambda$ of rank $r$ any Hamiltonian $\mathcal{C}(n,r-1)$-subcomplex $C\subset\mathcal{C}(P,\Lambda)$ inducing $\Lambda$ corresponds to a hyperelliptic involution $\tau_{C}\in\mathbb{Z}_{2}^{r}$. Our second main result (Theorem 4.40) is that in dimensions $n\leqslant 4$ this correspondence is a bijection.
@@
In Subsection 4.3 using the notion of a bipartite Hamiltonian subcomplex we reformulate (Corollary 4.20) the four color theorem for a simple $3$-polytope $P$ in terms of a disjoint union of simple edge-cycles containing all the vertices of $P$.
@@
  • In Proposition 4.29 and Corollary 4.30 we describe $k$-faces of a Hamiltonian subcomplex $C\subset\mathcal{C}(P,c)$ as unions of $k$-faces of $\mathcal{C}(P,c)$ separated by (k1)(k-1)-faces of defining faces.
@@
  • In Corollary 4.33 we describe explicitly the subgroup $H(C)\subset\mathbb{Z}_{2}^{m}$ corresponding to a $\mathcal{C}(n,k)$-subcomplex $C\subset\partial P$.
@@
  • In Construction 4.39 we show that any proper Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C\subset\mathcal{C}(P,c)$ induces a canonical vector-coloring $\widetilde{\Lambda}_{C}$ of rank $(k+1)$ such that $\mathcal{C}(P,c)=\mathcal{C}(P,\widetilde{\Lambda}_{C})$ and $N(P,\widetilde{\Lambda}_{C})$ is a hyperelliptic manifold with a canonical hyperelliptic involution in $\mathbb{Z}_{2}^{k+1}$ corresponding to $C$.
  • In Theorem 4.40 we formulate our second main result, namely that for $n\leqslant 4$ and a vector-coloring of rank $r$ of a simple $n$-polytope $P$ such that $N(P,\Lambda)$ is a closed manifold hyperelliptic involutions in $\mathbb{Z}_{2}^{r}$ are in bijection with proper Hamiltonian $\mathcal{C}(n,r-1)$-subcomplexes $C\subset\mathcal{C}(P,\Lambda)$ inducing $\Lambda$.
@@
In Subsection 4.7 we study geometric structures on hyperelliptic manifolds $N(P,\Lambda)$. Our main tool is Construction 4.46 of a geometric hyperelliptic manifold using a proper Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C$ in the boundary of a compact right-angled polytope $P$ in some geometry $\mathbb{X}$. This construction is a direct generalization of the construction from [M90, VM99M, VM99S2] based on a Hamiltonian cycle, theta- or $K_{4}$-subgraph in the boundary of a right-angled $3$-polytope.
@@
  • In Corollary 4.50 we show that $\mathbb{I}^{n}$ does not admit Hamiltonian $\mathcal{C}(n,k)$-subcomplexes for $n\leqslant 4$. In particular, the geometry $\mathbb{R}^{n}$ does not admit Construction 4.46 for $n\geqslant 4$.
@@
  • In Corollaries 4.52 and 4.54 and Remark 4.53 we show that if $P$ admits a Hamiltonian $\mathcal{C}(n,n-1)$-subcomplex, then each defining face is a polytope with only even-gonal 22-faces.
@@
  • In Proposition 4.57 we show that if a $4$-polytope $P$ admits a proper Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then $P$ has at least one $2$-face with three or four edges. In particular, the geometry $\mathbb{L}^{4}$ does not admit Construction 4.46, as well as $\mathbb{L}^{n}$ for $n\geqslant 5$ since in the latter geometries there are no compact right-angled polytopes due to results by V.V. Nikulin [N82] (see [V17]).
@@
  • In Proposition 4.58 we show that if a $4$-polytope $P=Q\times I$ admits a proper Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then one of the defining faces is a triangle. In particular, the geometries $\mathbb{L}^{3}\times\mathbb{R}$ and $\mathbb{L}^{2}\times\mathbb{R}^{2}$ do not admit Construction 4.46.
@@
  • In Proposition 4.59 we show that $\Delta^{n}$ admits a Hamiltonian $\mathcal{C}(n,n)$-subcomplex. In particular, the geometry $\mathbb{S}^{n}$ admits Construction 4.46 for $n\geqslant 2$.
@@
  • In Proposition 4.62 we show that the prism $\Delta^{n-1}\times I$ admits Hamiltonian $\mathcal{C}(n,n)$ and $\mathcal{C}(n,n+1)$-subcomplexes for $n\geqslant 4$. In particular, the geometry $\mathbb{S}^{n-1}\times\mathbb{R}$ admits Construction 4.46 for $n\geqslant 3$.
@@
  • In Construction 4.63 we describe the operation of cutting off a face and show that the initial polytope defines a proper Hamiltonian subcomplex in the resulting polytope. As a result in Corollary 4.69 we show that the polytope $P^{n}=\Delta^{p}\times\Delta^{q}$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex for $p,q\geqslant 1$. In particular, the geometry $\mathbb{S}^{p}\times\mathbb{S}^{q}$, $p,q\geqslant 2$, admits Construction 4.46.
@@
  • In Proposition 4.71 we show that if $P^{n}$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex such that its defining faces $\{M_{1},\dots,M_{s}\}$ do not intersect a facet $F_{i}\simeq\Delta^{n-1}$, then for any $k\geqslant 1$ the polytope $\Delta^{k}\times P$ has a Hamiltonian $\mathcal{C}(n+k,n+k+1)$-subcomplex, and in Corollary 4.72 we show that for any $p,q\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times\Delta^{q}\times I$ has a Hamiltonian $C(n,n+1)$-subcomplex. In particular, the geometries $\mathbb{S}^{p}\times\mathbb{S}^{q}\times\mathbb{R}$, $p,q\geqslant 2$, and $\mathbb{S}^{p}\times\mathbb{R}^{2}$, $p\geqslant 2$, admit Construction 4.46.
@@
  • In Proposition 4.73 we show that if an $n$-polytope $P$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex $C$ such that its defining faces $\{M_{1},\dots,M_{s}\}$ do not intersect an (n2)(n-2)-face $G\subset P$, then $G\simeq\Delta^{n-2}$, and $P$ has a Hamiltonian $\mathcal{C}(n,n)$-subcomplex. Using this fact in Corollary 4.74 we show that $\Delta^{2}\times\Delta^{n-2}$ has a Hamiltonian $\mathcal{C}(n,n)$-subcomplex for $n\geqslant 3$, and in Corollary 4.75 we show that for $p\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times P_{k}$, where $P_{k}$ is a $k$-gon, $k\geqslant 3$, admits Hamiltonian $\mathcal{C}(n,n)$- and $\mathcal{C}(n,n+1)$-subcomplexes. In particular, the geometries $\mathbb{S}^{p}\times\mathbb{S}^{2}$, $\mathbb{S}^{p}\times\mathbb{R}^{2}$ and $\mathbb{S}^{p}\times\mathbb{L}^{2}$, $p\geqslant 2$, admit Construction 4.46.
@@
  • In Example 4.76 we show that the polytope $P_{5}\times P_{5}$ admits a Hamiltonian $\mathcal{C}(4,5)$-subcomplex. In particular, the geometry $\mathbb{L}^{2}\times\mathbb{L}^{2}$ admits Construction 4.46.
@@ 2. Rational homology $4$-spheres among $N(P,\Lambda)$
the Poincare-Lefschetz duality $H_{k}(X,\partial X)\simeq H^{n-k}(X)$ and the universal coefficients formula. ∎
@@
[E24, Lemma 2.2] Let a point $p\in\partial P$ belong to exactly $l\geqslant 1$ facets $G_{i_{1}}$, $\dots$, $G_{i_{l}}$ of $\mathcal{C}(P,c)$. Then there is a piecewise linear homeomorphism $\varphi$ of a neighbourhood $U\subset P$ of $p$ such that $U\cap G_{j}=\varnothing$ for $j\notin\{i_{1},\dots,i_{l}\}$ to a neighbourhood $V\subset\mathbb{R}^{l}_{\geqslant}\times\mathbb{R}^{n-l}$ such that $\varphi(G_{j_{s}}\cap U)=V\cap\{y_{s}=0\}$, $s=1,\dots,l$.
@@
Let $P$ be a simple $n$-polytope, and let $\omega_{1},\omega_{2}\subset[m]$ be subsets such that $\omega_{1}\cap\omega_{2}=\varnothing$, and $P_{\omega_{1}}$, $P_{\omega_{2}}$ are (n1)(n-1)-RHDRHD.
@@
Let $P$ be a simple $n$-polytope, $n\leqslant 4$, and let $\omega_{1},\omega_{2}\subset[m]$ be subsets such that $\omega_{1}\cap\omega_{2}=\varnothing$, and $P_{\omega_{1}}$, $P_{\omega_{2}}$ are (n1)(n-1)-disks.